# Dinkel: Testing Graph Database Engines via State-Aware Query Generation

### v2

@@ 3. State-Aware GDBMS Testing
3.1. Graph State Modeling
Cypher query execution can be affected by two kinds of graph state, which we refer to as query context and graph schema, respectively.
@@ 3.1. Graph State Modeling
Query Context. One kind of graph state information is the temporary variables declared in the query. These variables have specific types and scopes, affecting only the query where they are declared. They can be either concrete values with specific types (e.g., INTEGER 0) or aliased to specific nodes or relationships (e.g., node x and relationships y in Figure 3). Such variables can be referenced only after they are defined and within their scope. We refer to these variables, their types, and their scope information as query context. Query context may change at different Cypher clauses. Specifically, query context includes new information when a new variable is declared by a clause, and excludes outdated information when the clause is out of the scope of existing variables.
Graph Schema. Another graph state information is the schema of the stored graph data. It includes graph labels and properties (e.g., label A and B for relationships in Figure 3). We refer to such information as graph schema. Graph schemas can be changed by specific Cypher clauses. For example, CREATE clauses can create nodes or relationships with new labels and properties, while REMOVE clauses can remove existing labels or properties.
@@
Different from query context, whose effects are limited by their scope, the operations (e.g., CREATE clause that declares new labels) on graph schema affect the manipulated database permanently. In addition, graph schema can be referenced even though the labels and properties are non-existent in the manipulated graph. This design improves query flexibility as users can write valid queries without concerning the current graph schema.
1CALL {
2 WITH 0 AS x
3 CREATE ()-[:A {n0:x}]->()
4}
5MATCH ()-[:(B & C) {n0: 0}]->(:!(D & E))
6RETURN 0
Figure 6. A Cypher query triggering a Neo4j internal error: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.type()” because ”this.v3_relationships” is null).
Table 2. State changes of the query in Figure 6
Line Query Context Graph Schema
Label Property
1 {} {} {}
2 {x, INT, CALL} {} {}
3 {x, INT, CALL} {A} {n0}
4 {} {A} {n0}
5 {} {A} {n0}
6 {} {A} {n0}
Example. To better understand query context and graph schema, we discuss the Cypher query shown in Figure 6, which causes an internal error in the enterprise version of Neo4j, which is close-source. This query involves complex graph state changes by using multiple Cypher clauses. Table 2 shows the concrete state at each line of the query. At the first line of the query, query context and graph schema contain no state information as the CALL clause has not had any effects yet. At the second line, a WITH clause declares a temporary variable with value 0, and thus the query context now contains the corresponding variable identifier (i.e., x), its type, and its scope. Because the WITH clause is inside the CALL clause, the scope of the variable x is limited to the CALL clause. At the third line, the CREATE clause updates the graph schema by adding a new relationship with label A and properties n0 referencing variable x. The new graph schema (i.e., label A and property n0) is not limited by the scope of the CALL clause and still exists after the CALL clause and even when the query finishes. At the fifth line, the MATCH clause references the property n0 and four non-existing labels B, C, D, and E. The query ends with a RETURN clause.
Generating such a query is challenging because it involves multiple clauses (5 clauses) and graph data references (e.g, the CREATE references variable x, and the MATCH references property n0). Utilizing the two abstractions, query context and graph schema, our approach can explicitly and precisely describe the graph state maintained by GDBMSs during query processing, which enables our approach to accurately reference available graph data in the generated queries.
@@ 3 Function GenQuery():
7             AppendQuery(query, clause);
@@
       // initialize the clause according to its type
@@
17       foreach subclause, subtype in clause do
18             subclause, qc, gs $\leftarrow$ GenClause(qc, gs, subtype);
@@ 20       CleanQC(qc);
3.2. State-Aware Query Generation
Insight. Our state-aware query generation is based on an observation that both query context and graph schema are updated only when Cypher clauses are invoked (e.g., CREATE and MATCH in Figure 6), or exit specific clause contexts (e.g., CALL in Figure 6). Utilizing this observation, our approach can accurately track the state changes inside the Cypher queries by analyzing the possible impact caused by specific clauses in the queries.
@@ 3.2. State-Aware Query Generation
Algorithm. Algorithm 1 shows the procedure of our state-aware query generation. Our approach does not need any input. Initially, it constructs an empty query that contains no clause, and initializes query context qc and graph schema gs to empty sets (line 1-1). Then, the approach incrementally appends the query with a newly generated clause (line 1-1). Each time a new clause is generated, the query context and graph schema may be updated (line 1). Our approach decides whether to append the query with more clauses based on a certain probability P. If the length of the query reaches the limit Lmax, the approach stops appending the query (line 1). In the end, the generated query is returned (line 1).
@@
To generate a clause, our approach needs to reference the query context and graph schema. If the clause type is not specified, the approach randomly chooses a clause type (line 1-1). According to the elements in query context, graph schema, and the specified clause type, our approach randomly initializes a clause (line 1). Initialization for different clauses can vary (e.g., the procedures for CALL and CREATE are different), but it generally includes the procedure of determining the number and types of the components needed by the clause, generating corresponding expressions for the components, and analyzing the impacts of the generated clause. If the generated clause contains subclauses, the approach will recursively call GenClause() to generate subclauses with specified types (line 1-1). After the clause is generated, the approach cleans up the query context if necessary (line 1), such as removing out-of-scope variables at the end of a CALL clause. In the end, the generated clause, the updated query context, and the graph schema are returned (line 1).
Figure 7. Generation process for the query in Figure 6.
@@
Example. Figure 7 shows how our approach generates the query in Figure 6. Initially, the query is empty, and query context and graph schema contain nothing (step ➀). The approach randomly determines that the first clause to be generated is a CALL clause (step ➁). When initializing the CALL clause, our approach determines its components, which are two subclauses. Then, these two subclauses are recursively generated. The first subclause is randomly constructed as a WITH clause with two components, 0 and x. After the WITH subclause is completed, the approach updates the query context by adding a temporary variable x with its type and scope, according to the functionality of WITH (Cypher Query Language Reference., 2022) (step ➂). The approach continues to generate the second subclause, which is randomly determined as a CREATE clause. The component of the CREATE clause is generated with graph pattern ()-[:A {n0:x}]->(), which references the existing variable x. After the CREATE is constructed, the approach updates the graph schema with the newly generated label A and property n0 (step ➃). After the two subclauses are constructed, the CALL generation is completed. Before generating other clauses, our approach cleans the variable x in the query context, because it is scoped to only the CALL clause (step ➄). Then, our approach continues to generate a MATCH clause, which references the existing property n0, according to the latest graph schema, and four non-existing and randomly constructed labels B, C, D, and E (step ➅). In the end, our approach generates a RETURN clause to finish the query (step ➆).
Table 3. Effects of Cypher clauses on graph states
Clause Query Context Graph Schema
Label Property
MATCH + / /
CREATE + + +
MERGE + + +
DELETE / / /
REMOVE / - -
SET / +/- +/-
UNWIND + / /
WITH + / /
RETURN + / /
FOREACH +/- +/- +/-
CALL +/- +/- +/-
UNION +/- +/- +/-
EXISTS +/- +/- +/-
COUNT +/- +/- +/-
  • /: the clause has no effect on the graph state;
    +: the clause can add elements to the graph state;
    -: the clause can remove elements from the graph state;
    +/-: the clause can both add and remove elements.
Clause Impact Analysis. Different Cypher clauses have different effects on query context and graph schema, and our approach models their effects and supports the clause generation accordingly (e.g., RandInitClause function in Algorithm 1). Table 3 lists the effects of different clauses, according to the openCypher specification (Cypher Query Language Reference., 2022). Most clauses can modify either query context, graph schema, or both. The DELETE clause does not affect either query context or graph schemas, because it manipulates only data but does not affect available references or graph attributes. For example, in Figure 3, variable y can still be referenced despite being operated on by a subsequent DELETE clause. Five clauses (i.e., FOREACH, CALL, UNION, EXISTS, and COUNT) can involve subqueries using subclauses, and they can have effects on both query context and graph schemas via invoking specific subclauses. In addition, these clauses demarcate the scope of the query contexts defined by the subclauses. For example, the CALL clause in Figure 6 modifies both query context and the graph schema via the WITH and CREATE clauses, and the variable x defined by the WITH clause can be referenced only within the CALL clause.
@@ 4. Implementation
We implement our approach into a practical and fully automatic GDBMS testing framework, Dinkel, integrating Cypher query generation, bug-triggering query reduction, and efficient bug deduplication. Figure 8 shows its architecture. The overall code base of Dinkel consists of 10k lines of Go code. The source code of Dinkel will be available on GitHub. The following discusses important implementation details.
@@
Supported Clauses in Query Generation. As of writing this paper, Dinkel supports all non-administrative clauses (e.g., all the clauses shown in Table 3) in the tested GDBMSs. For example, in Neo4j, 21 out of the total 25 clauses are non-administrative and thereby supported. Administrative clauses, such as SHOW FUNCTIONS and USE, are not supported because they are not related to the logic of query processing in GDBMSs. For example, The clause SHOW FUNCTIONS is designed to list all available functions and cannot be combined with other clauses. The clause USE is used to switch the manipulated databases, which is not necessary as all databases are initialized to empty and they function in the same way.
@@
Compared to existing approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024), Dinkel supports more Cypher clauses, as shown in Table 4. None of the existing approaches supports FOREACH clauses, because this clause complicates the Cypher control flows and can significantly change graph states during query execution. UNION, EXISTS, and COUNT are not supported as well, as these clauses invoke subqueries that can inherit and change the graph states of the main query. Lacking a systematic model for handling graph state changes, existing approaches are limited in supporting various Cypher clauses.
Bug Detection. In this paper, we focus on the query generation of GDBMS testing and apply only general methods for identifying triggered bugs. Specifically, for each executed query, Dinkel checks the messages returned from the GDBMS. If the message indicates an internal error in the GDBMS (e.g., the error in Figure 6), a bug is identified. We also enable all the assertions embedded in the tested GDBMSs to catch assertion failures. In addition, we enable ASan (Address Sanitizer., 2019) for the GDBMSs developed in C/C++ to identify memory bugs triggered by our generated queries.
Figure 8. Achitecture of Dinkel.
Query Reduction. No existing tool is available for reducing Cypher queries, which makes it difficult for developers to minimize bug-triggering queries and investigate bugs. To address this problem, we implemented Dinkel with an automatic query reduction method. The core idea of this method is to reduce queries clause by clause. For each Cypher clause, Dinkel first tries to delete it. If the query without the clause still triggers the bug, the clause will be permanently removed. Otherwise, the deleted clause will be recovered, and Dinkel goes on to try to replace the clause with an alternative clause if possible. For example, Dinkel can try to replace the CALL clause in Figure 6 with its subclause CREATE. If some clauses are successfully deleted or replaced in one try, Dinkel will restart the reduction process for the reduced queries. The process stops when no clause in the query can be further reduced. With this method, bug-triggering queries can be efficiently reduced.
Bug Deduplication. When a query generated by Dinkel triggers a bug, we need to demonstrate whether this bug is unique or a duplicate of another bug already found by Dinkel. A feasible way is bisection (Git-bisect: Use binary search to find the commit that introduced a bug., 2024; Zeller, 2009), which finds the earliest bug-inducing commits for each triggered bug and deduplicates bugs by checking whether their bug-inducing commits are the same. As GDBMSs are large-scale system software, compiling and building a version of a GDBMS demands considerable time and resources (e.g., 8 minutes for Neo4j in our evaluation), which can make bisection infeasible when a large number of commits need to be bisected (e.g., 79K commits in Neo4j’s GitHub repository (Neo4j. Graphs for Everyone., 2024)). To improve the efficiency of bisections, Dinkel leverages containerization techniques (e.g., Docker (Docker: Accelerated Container Application Development., 2024)) to cache and manage the built GDBMS versions. Given a batch of bugs needed to be deduplicated, Dinkel sets up multiple containers to bisect the bug-inducing commits. During bisection for a bug, Dinkel needs to test many commits by checking whether the bug can be reproduced in this commit version. For each commit that has not been tested yet, Dinkel compiles and builds a GDBMS version, and caches the built version in a container image. When testing a cached commit during bisection for a new bug, Dinkel directly utilizes the image to set up the container, without repeatedly compiling and building the same versions of GDBMSs.

### v3

@@ 3.1. State-Aware Query Generation
3.1.1. Graph State Modeling.
Cypher query execution can be affected by two kinds of graph states:
@@
Query Context. One kind of graph state are the temporary variables declared in the query. These variables have specific types and scopes, affecting only the query state where they are declared. They can either be concrete values with specific types (e.g., integer 0) or aliased to specific nodes or relationships (e.g., node x and relationship y in Figure 1). Such variables can be referenced only after they are defined, within their scope. We refer to these variables, their types, and their scope information as query context. Query context may change at different Cypher clauses. Specifically, query context includes new information when a new variable is declared by a clause, and excludes outdated information when the clause is out of the scope of existing variables.
@@
Graph Schema. Another graph state is the schema of the stored graph data. It includes graph labels and properties (e.g., label A and B for relationships in Figure 1). We refer to such state information as graph schema. Graph schemas can be changed by specific clauses. For example, CREATE clauses can create nodes or relationships with new labels and properties, while REMOVE clauses can remove existing labels or properties. Different from query context, whose effects are limited by their scope, the operations (e.g., CREATE clause that declares new labels) on graph schema affect the database permanently. In addition, graph schema can be referenced even though the labels and properties are nonexistent. This design improves query flexibility as users can write valid queries without concerning themselves with the current graph schema.
3.1.2. Query Generation
Our generation is based on a key insight that clauses are the basic units for operating on graphs, and both query context and graph schema are updated only when clauses are invoked or exit specific clause context. Based on this insight, our approach track the state changes inside Cypher queries by analyzing the possible impact caused by specific clauses in queries.
@@ 3 Function GenQuery():
7    AppendClause(query, clause);
@@
 // initialize the clause accordingly
@@
17 foreach subclause, subclause_type in clause do
18    subclause, qc, gs $\leftarrow$ GenClause(qc, gs, subclause_type);
@@ 21 CleanQC(qc);
Algorithm. Algorithm 1 shows the procedure of our state-aware query generation. Our approach does not need any input. Initially, it constructs an empty query that contains no clause, and initializes query context qc and graph schema gs to empty sets (line 1-1). Then, the approach incrementally appends the query with a newly generated clause (line 1-1). Each time a new clause is generated, the query context and graph schema may be updated (line 1). Our approach decides whether to append the query with more clauses based on a certain probability P. If the length of the query reaches the limit Lmax, the approach stops appending to the query (line 1). In the end, the generated query is returned (line 1).
@@
To generate a clause, our approach references the query context and graph schema. If the clause type is not specified, the approach randomly chooses a clause type (line 1-1). According to the elements in query context, graph schema, and the specified clause type, our approach randomly initializes a clause (line 1). Initialization for different clauses can vary (e.g., the procedures for CALL and CREATE are different), but it generally includes the procedure of determining the number of components needed by the clause and their types, generating corresponding expressions for these components, and analyzing the impacts of the generated clause. If the generated clause contains subclauses, the approach will recursively call GenClause() to generate subclauses with specified types (line 1-1). After the clause is generated, the approach cleans up the query context if necessary (line 1), such as removing out-of-scope variables at the end of a CALL clause. In the end, the generated clause, the updated query context, and the graph schema are returned (line 1).
Figure 5. Generation for the original query in Figure 2.
3.1.3. Illustrative Example
@@
We show the effectiveness of our generation using a complex query example, the original query in Figure 2. As shown in Figure 5, initially, the query is empty, and both query context and graph schema contain nothing (step ➀). The approach randomly determines that the first clause is a RETURN clause with an integer 0 as the returned value n1. Then, the approach updates the query context by adding n1 with its type and scope (step ➁). The second clause is a UNION clause. After constructing the UNION, the approach cleans the query context as the clause after the UNION cannot reference the variables defined before the UNION (step ➂). Subsequently, the approach randomly decides to generate a CALL clause with a subclause (step ➃). The subclause is recursively generated as a FOREACH clause, defining n2 in its loop condition. The FOREACH also has a subclause, which is another FOREACH clause, defining n3. The last FOREACH has a MERGE as its subclause. Our approach records the query context (i.e., n2, n3 defined by the two FOREACH clauses) and graph schema (i.e., label A and property u referenced by the MERGE) updated in this recursive generation process (step ➄). After completing generation of the FOREACH clauses, the approach removes n2 and n3 from query context as they are scoped to only the FOREACH clauses. Finishing the CALL, our approach continues to randomly generate clauses, including a MATCH defining y, and a RETURN referencing y as n1, and update the query context by adding y and n1 (step ➅). Finally, our approach randomly generates a UNION followed by a CREATE and a RETURN. The UNION cleans all the contents of the query context, after which RETURN adds a new variable n1 (step ➆).
3.2. Fine-Grained Query Transformation
To enable general query validation, we construct test oracles based on the fine-grained units of queries: clauses and expressions. Each query consists of a chain of clauses, which can be transformed by modifying the clauses or prepending them with new clauses. Each clause may include expressions, on which more transformations can be applied by using algebraic identities or Cypher features.
Table 2. Clause-level transformation rules. (1) T.S.: transformation strategies including Dead Code (DC), Unused Context (UC), Redundant Write (RW), and Inconsequential Supersession (IS); (2) <literal>: any atomic expression; (3) <mirror(x)>: a mirrored version of x (e.g., a->b $\Rightarrow$ b<-a); (4) <clauses>: a random number of random clauses; (5) <new pattern>: a graph pattern that involves at least one previously unused label or property
Type Original T.S. Transformed
Read- Only C DC CALL {MATCH x WHERE false <clauses>} C
DC CALL {MATCH <new pattern> <clauses>} C
IS OPTIONAL MATCH <new pattern> C
UC CALL {WITH * RETURN x AS i} C
UC UNWIND <literal> AS i C
WITH x UC WITH x, y AS i
MATCH x IS MATCH x WHERE true
IS MATCH <mirror(x)>
Write/ Project C DC FOREACH (i IN [] |<clauses>) C
RW DELETE <deleted var> C
RW CREATE x DELETE x C
RW SET x += {} C
REMOVE x RW REMOVE x, x
DELETE x RW DELETE x, x
IS DETACH DELETE x
Any C IS WITH * C
3.2.1. Clause-Level Transformations
These transformations manipulate queries by either adding new clauses, or switching out clauses with equivalent counterparts. Table 3.2 lists the transformation rules. We design different rules for reading clauses and writing/projecting clauses because the Cypher specification limits the kinds of clauses that can be used before specific clauses. The transformations consider state information to avoid semantic invalidity (e.g., repeatedly defining the same variables).
We introduce four strategies to achieve semantic-preserving transformation for clauses: Dead Code (DC), Unused Context (UC), Redundant Write (RW), and Inconsequential Supersession (IS). Dead Code expands the target clause by introducing additional clauses that, due to Cypher’s way of handling queries, will not be executed. We implement three rules for this strategy. Two rules transform read clauses by adding a CALL clause with a MATCH. The MATCH clause is designed to return no rows, leading to no rows for the clauses following the MATCH to ingest. As a result, the subsequent clauses in the CALL clauses will not be executed. One rule transforms write clauses by inserting a FOREACH before them. The FOREACH iterates the elements in a given array and invokes clauses to operate on the iterated elements. Because the array is empty, no iterations are performed by the FOREACH. Therefore, the invoked clauses in the FOREACH will not be executed. For example, in Figure 2, our approach transforms the first RETURN clause of the original query by adding a dead FOREACH, which preserves the semantics of the RETURN and thus should not affect the final results. However, the results of the transformed query change, indicating that a bug was triggered.
Unused Context expands a clause by involving unused query context. Specifically, it defines new variables that are not referenced by subsequent parts of the query. Redundant Write is straightforward. It expands a clause by deleting a previously deleted node, removing the same property multiple times, or creating and immediately deleting a subgraph. Lastly, Inconsequential Supersession expands a clause by adding inconsequential operations (e.g., adding WHERE true to a MATCH clause) or modifying the formats of the clauses (e.g., reversing the graph pattern of a MATCH clause).
Clause-level transformations have the ability to influence the query context through manipulating variables, as well as changing the database state by modifying graph elements or affecting control flow. By doing so, Dinkel can effectively force the GDBMSs to execute system logic that was not covered by the original queries, and thus validate the query results by checking whether the different executed logic produces the same results.
Table 3. Expression-level transformation rules. (1) T.S.: transformation strategies including Mathematical Identity (MI) and Cypher Feature (CF); (2) P': randomly generated boolean expression; (3) E': randomly generated expression
Type Initial T.S. Transformed
Bool P MI NOT (NOT P)
MI P OR false
MI P AND true
P OR Q MI NOT ((NOT P) AND (NOT Q))
P AND Q MI NOT ((NOT P) OR (NOT Q))
true MI true OR P'
MI P' OR (NOT P') OR (P' IS NULL)
false MI false AND P'
MI P' AND (NOT P') AND (P' IS NOT NULL)
NULL MI NULL {XOR|=|<>|<|>|>=|<=} P'
Num x MI -(-x)
MI x {+|-} 0
MI x {*|/|^} 1
0 MI 0 * x
NULL MI NULL {+|-|*|/} x
Str S MI S + ””
List L MI L + []
Any E IS NULL MI NOT (E IS NOT NULL)
E CF CASE WHEN true THEN E ELSE E' END
CF CASE WHEN false THEN E' ELSE E END
3.2.2. Expression-Level Transformations
These transformations operate on expressions within Cypher clauses. Although they do not modify the state of the database or query, they still force the GDBMS to evaluate more complex expressions. Table 3.2.1 lists our expression-level transformations.
Our expression-level transformations involve two strategies: Mathematical Identity (MI) and Cypher Feature (CF). Mathematical Identity leverages mathematical laws (e.g., De Morgan’s laws for boolean con- and disjunctions, and double negations) to transform the expressions while preserving their semantics. Cypher Feature leverages the features of the Cypher language. For example, Dinkel can transform an arbitrary expression into a CASE expression, where the branch condition is determined to be true, and the corresponding value of the true branch is the original expression. When transforming expressions, Dinkel sometimes needs to involve new components in the transformed expressions (e.g., the E' in the transformed CASE expressions). In this case, Dinkel constructs these components by referencing the available variables recorded in the corresponding query context.
@@ 4. Implementation
We implement our approach as a fully automatic GDBMS testing framework, Dinkel. The overall codebase of Dinkel consists of 10k lines of Go code. The source code of Dinkel is available for review on an anonymized repository [^1] https://github.com/Anon10214/dinkel and will be public concurrent with the publication of the paper. The following discusses important implementation details.
@@
Supported Clauses in Query Generation. Dinkel supports all non-administrative clauses in the tested GDBMSs. For example, in Neo4j, 21 out of the total 25 clauses are non-administrative and thereby supported. Administrative clauses, such as SHOW FUNCTIONS and USE, are not supported because they are not related to query processing logic. For example, SHOW FUNCTIONS is designed to list all available functions and cannot be combined with other clauses. USE is used to switch the manipulated databases, which is not necessary as all databases are empty on initialization and they function in the same way.
@@
As shown in Table 4, Dinkel supports more Cypher clauses than existing approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024b). None of the existing approaches support FOREACH clauses, because this clause complicates control flow and can significantly modify graph states during query execution. UNION, EXISTS, and COUNT are not supported either, as these clauses invoke subqueries that can inherit and change the graph states of the main query. Lacking a systematic model for handling graph state changes, existing approaches are limited in generating complex queries using these clauses. Moreover, existing approaches support validation for only queries following specific patterns, while many clauses complicate the semantics of Cypher queries, making them unable to be validated by existing approaches.
@@
Bug Detection. Dinkel detects logic bugs via its fine-grained query transformation. Besides logic bugs, Dinkel applies some methods to identify internal errors and memory bugs in GDBMSs. Specifically, for each executed query, Dinkel checks the messages returned from the GDBMS. If the message indicates an internal error in the GDBMS (e.g., the error in Figure 6), a bug is identified. We also enable all embedded assertions in the tested GDBMSs to catch assertion failures. In addition, we enable ASan (Address Sanitizer., 2019) for the GDBMSs developed in C/C++ to identify memory bugs triggered by our generated queries.
Query Reduction. No existing tool is available for reducing Cypher queries, which makes it difficult for developers to minimize bug-triggering queries and investigate bugs. To address this problem, we implemented Dinkel with an automatic query reduction method. The core idea of this method is to reduce queries clause by clause. For each clause, Dinkel tries to delete it. If the query without the clause still triggers the bug, the clause will be permanently removed. Otherwise, the deleted clause will be recovered, and Dinkel goes on to try to replace the clause with an alternative clause if possible. For example, Dinkel can try to replace the CALL clause in Figure 2 with its subclause CREATE. If some clauses are successfully deleted or replaced in one try, Dinkel will restart the reduction process for the reduced queries. The process stops when no clause in the query can be further reduced.